Hi I would like some help with this equation, I get confused because of the double d.I have tried opposite in each side but I don't think is right.I appreciate it.
6d+7=5d-2

Found 2 solutions by rfer, derangoj:
Answer by rfer(16322)

6d-5d=-2-7
d=-9

Answer by derangoj(14)

6d+7=5d-2
subtract 5d-2 from both sides you get.
6d+7
-5d+2
d+9=0
d=-9